# Modulation of Polarization and Metallicity in Janus Sliding Ferroelectrics

### v1

@@ Introduction
Figure 6(a) and Figure 6(b) show how the polarization obtained from charge density ($P_{CD}$) and the electronic bandgap($E_{gap}$) vary with $d$ for the XMoY bilayers. As $d$ decreases, $P_{CD}$ increases, whereas $E_{gap}$ decreases. Notably, the TeMoS bilayer becomes semi-metallic when $d$ is reduced by more than 4%, as illustrated in Figure 6(b) and Figure S11 [33]. A comparison of the percentage changes in $P_{CD}$ and $E_{gap}$ with $d$, shown in Figure S12 [33], reveals that E-in stacked Janus bilayers exhibit a greater variation in $E_{gap}$ and a smaller variation in $P_{CD}$ compared to their E-out stacked counterparts.
@@
Figure 6(c) and Figure 6(d) display the variation of the VBM and CBM ratios, respectively, with $d$ for SeMoS (E-in stacking) and SMoSe (E-out stacking) bilayers. Both ratios show a slight overall increase with increasing $d$, with more pronounced changes in the CBM ratios. These variations are generally monotonic, except at points where the VBM or CBM undergoes a band-edge shift. For example, a shift in the VBM is observed at a 5% increase in $d$ for the SMoSe bilayer (Figure 6(c)), and a shift in the CBM occurs at a 5% decrease in $d$ for the SeMoS bilayer (Figure 6(d)). Figure S13 and Figure S14[33] present similar trends in the VBM and CBM ratios across all XMoY bilayers, further confirming the correlation between interlayer distance and band-edge contributions.
@@
As shown in Figure S13 (c) [33], a notable shift of the VBM from the K point to the $\Gamma$ point occurs in the MoTe2 bilayer with just a 1% decrease in $d$. This shift leads to a sudden increase in the VBM ratio, indicating a reduced depolarization effect from hole doping when $d$ is decreased beyond 1%. Interestingly, a similar behavior is observed for MoTe2, WSe2, and SWSe bilayers when SOC is turned off: the VBM shifts and the VBM ratio increases, as shown in Figure S8 [33], which results in a reduction in hole-doping-induced depolarization, as shown in Figure S9-S10 [33]. These observations suggest that decreasing the interlayer distance could likewise enhance the VBM ratio and reduce hole-doping-induced depolarization in WSe2 and SWSe bilayers.
@@
To validate our intuition regarding the enhancement of the VBM ratio with reduced interlayer distance ($d$) in WSe2 and SWSe bilayers, we analyze the variation of VBM and CBM ratios with $d$, as shown in Figure 6 (e)-(f). Both WSe2 and SWSe bilayers exhibit a sudden increase in the VBM ratio when $d$ is decreased beyond 3%, which corresponds to a shift of the VBM from the K point to the $\Gamma$ point, as shown in Figure 6(e). In contrast, the CBM ratio increases monotonically with increasing $d$, as shown in Figure 6 (f), similar to the behavior observed in the XMoY bilayers. The enhanced VBM ratios for WSe2 and SWSe bilayers closely resemble those obtained when the SOC effect is switched off, as shown in Figure S8 [33]. Therefore, we anticipate a comparable reduction in hole-doping-induced depolarization in WSe2 and SWSe bilayers for reduced $d$, consistent with the effect observed upon disabling SOC (Figure S10 [33]).
@@
Motivated by the potential for enhanced polarization with multiple stable polarization states [25, 29], and the possibility of increased interlayer vdW interactions when only one Janus monolayer is present at the interface [36, 37], we propose an ABC-stacked trilayer structure. In this design, a MoS2 monolayer is sandwiched between two Janus monolayers, with the intrinsic electric fields ($E$) in the Janus layers oriented toward the heterostructure interface, as illustrated in Figure 7(a).
The ABC stacking configuration shown in Figure 7(a) closely resembles the AB stacking illustrated in Figure 1(a), producing an upward out-of-plane polarization at each interface. The interlayer distances at the top and bottom interfaces, denoted as $d_{1}$ and $d_{2}$, respectively, are key factors in modulating the vdW interactions and the resulting interfacial polarizations. As shown in Table S4 [33], $d_{1}$ and $d_{2}$ have nearly identical values, indicating similar interlayer interactions and, consequently, comparable polarization at both interfaces, which is consistent with the behavior expected from interfacial ferroelectrics.
@@
The trilayer systems have lower (more negative) formation energies than their corresponding bilayer systems, indicating enhanced stability, as shown in Figure 7(b). We further validated the dynamical stability of both the Janus trilayers and their bilayer counterparts by calculating their phonon dispersion curves (Figure S16 [33]), which show no imaginary frequencies.
@@
In Figure 7(c), we compare the polarization values for the trilayer systems with those of their bilayer counterparts. Interestingly, similar to the bilayers, the trilayer Janus sliding ferroelectrics exhibit higher polarization when the $E$ fields point toward the interface. These $E$ fields also lead to a reduction in the interlayer distances, $d_{1}$ and $d_{2}$, as shown in Table S4 [33]. Additionally, the charge density profiles of the Janus trilayers, shown in Figure S17 [33], display similar modulation due to the $E$ fields as observed in the E-in stacked Janus bilayers. These results suggest that, in ABC-stacked Janus trialyers, the intrinsic $E$ fields modulate the polarization by redistributing interfacial electron density and tuning the interlayer spacing—mirroring the mechanism found in AB-stacked Janus bilayers.
The potential profiles in Figure S17 [33] confirm the interfacial nature of the polarization in the trilayer systems, which results in the cumulative nature of polarization at each interface. This explains the higher polarization values observed in Figure 7(c) compared to their bilayer counterparts. Additionally, the $\Delta\phi$ and $P_{CD}$ values follow a similar trend, further supporting the observed polarization behavior, as shown in Table S4 [33].
@@
The band structures of the trilayer systems exhibit strong asymmetry in the layer-wise contributions to the VBM and CBM. This leads to an uneven distribution of the partial charge density associated with the VBM and CBM, as illustrated in Figure S18 [33], with the asymmetry becoming more pronounced in the Janus sliding ferroelectric configurations. Consequently, based on the earlier discussion of doping-induced depolarization in Janus bilayers, we anticipate a significantly stronger depolarization effect from charge doping in trilayer systems compared to their bilayer counterparts.
@@
Finally, we would like to suggest some interesting future avenues for experimental and theoretical work motivated by our results. Experiments confirming our theoretical finding of enhancement of polarization along with bandgap and hole-doping-induced depolarization reduction by out-of-plane compressive strain for AB-stacked WSe2 bilayer would be an interesting and useful finding, considering the already existing experimental work on the doping-induced polarization modulation for WSe2 bilayer in Ref. [25]. Also, the semi-metallic nature observed after reducing the interlayer distance for the TeMoS bilayer suggests the possibility of designing Janus sliding ferroelectrics with intrinsic semi-metallicity by stacking them in E-in stacking, given much lower bandgaps for parent TMD bilayers. We expect this work to motivate several other theoretical and experimental investigations to design better-performing sliding ferroelectrics with additional properties such as a low bandgap and polar metallicity.
Computational Methods
Our first-principles calculations were carried out within the density functional theory (DFT) formalism, as implemented in the quantum espresso code [38, 39]. We used ultrasoft pseudopotentials [40] with a kinetic energy cut-off of 90 Ry for the wavefunctions and 720 Ry for charge density, with a $k$-mesh of $14\times 14\times 1$. We employed the Perdew-Burke-Ernzerhof (PBE) formalism to describe the exchange-correlation functional [41], along with the Grimme-D3 dispersion correction with BJ damping [42]. We used a vacuum of around 30Åto avoid interaction between adjacent images due to the periodic boundary conditions. The electronic self-consistent cycle used the convergence criterion of $1\times 10^{-8}$ eV. We relaxed the structures until the residual forces on the atoms were less than 0.001 eV/Å. We calculated the electric polarization values using the Berry phase method [34, 35], unless specified otherwise. The electrostatic potential profiles are obtained using the ionic and Hartree contributions from the charge density [24, 43]. We also used a dipole moment correction along the out-of-plane direction for the electrostatic potential profiles.
We implemented charge doping calculations using the fractional nuclear charge pseudo atom approach [44]. Following a similar approach as Ref. [25], we carried out the doping calculations by charging the metal nuclei. The charging of the metal nuclei was done by adding an equal amount of fractional charge to the metal nuclei of both layers for all the bilayer systems. The maximum doping limit for electron and hole dopants we used in this study is $\sim$1013 cm-2, which is the highest reported experimentally accessible charge density in such systems [25]. At each doping stage, we relaxed the bilayers keeping the lattice constants fixed. We used Fermi-Dirac smearing for the band structure calculations and for the self-consistent cycle during the charge doping to enhance the convergence. We employed spin-orbit interactions for band structure and charge doping calculations, as well as other cases as explicitly mentioned in the results section. We carried out phonon calculations within the finite displacement approach using a $2\times 2\times 1$ supercell, as implemented using a combination of quantum espresso and phonopy codes [45, 46].
Supporting Information
@@
Formation energy plot for bilayers; charge density and potential profiles for AB and BA stacking of SeMoS bilayer; charge density and potential profiles for XMoY bilayers; potential profiles for XMoY bilayers; correlation plots for $\Delta\phi$ and $P_{CD}$ with $P_{berry}$; additional results for XMoY bilayers with $d=d_{avg}$; band structures of XMoY bilayers; comparison of band structure of MoTe2, WSe2 and SWSe bilayer without and with SOC; effect of doping on polarization for bilayers; band structures for TeMoS bilayer showing semi-metallic nature at reduced interlayer distance; comparison of percentage change in bandgap and polarization with change in $d$ for XMoY bilayers; variation of the VBM and CBM ratios with change in $d$ for XMoY bilayers; additional results for change in $d$ for WSe2 and SWSe bilayers; phonon dispersion curves; charge density and potential profiles for trilayers; band structure with layer-wise contribution analysis for trilayers; tables for data for bilayer and trilayer systems presented and referred in the paper; methodology for $P_{CD}$ calculation (PDF).

### v2

@@ Introduction
Figure 6(a) and Figure 6(b) show how the polarization obtained from charge density ($P_{CD}$) and the electronic bandgap($E_{gap}$) vary with $d$ for the XMoY bilayers. As $d$ decreases, $P_{CD}$ increases, whereas $E_{gap}$ decreases. The variations in $P_{CD}$ and $E_{gap}$ with $d$ are also observed in parent TMD bilayers, further underscoring the crucial role of interlayer vdW interactions in tuning the properties of TMD-based sliding ferroelectrics. Notably, the TeMoS bilayer becomes semi-metallic when $d$ is reduced by more than 4%, as illustrated in Figure 6(b) and Figure S13 [SuppMat]. A comparison of the percentage changes in $P_{CD}$ and $E_{gap}$ with $d$, shown in Figure S14 [SuppMat], reveals that E-in stacked Janus bilayers exhibit a greater variation in $E_{gap}$ and a smaller variation in $P_{CD}$ compared to their E-out stacked counterparts.
To further investigate pressure-induced phase transitions and tunability, we compressed the XMoY bilayers up to a 10% reduction in $d$, with the results shown in Figure S13 [SuppMat]. As expected, additional compression led to further bandgap reduction accompanied by enhanced polarization, while the TeMoS bilayer remained semimetallic with a zero bandgap. Notably, at 8% and 9% reduction in $d$, both TeMoSe and MoTe2 also displayed zero bandgaps and semimetallic behavior, as confirmed by their band structures. These observations reinforce the possibility of pressure-induced transitions from semiconducting to semimetallic or metallic states in TMD bilayers.
@@
Figure 6(c) and Figure 6(d) display the variation of the VBM and CBM ratios, respectively, with $d$ for SeMoS (E-in stacking) and SMoSe (E-out stacking) bilayers. Both ratios show a slight overall increase with increasing $d$, with more pronounced changes in the CBM ratios. These variations are generally monotonic, except at points where the VBM or CBM undergoes a band-edge shift. For example, a shift in the VBM is observed at a 5% increase in $d$ for the SMoSe bilayer (Figure 6(c)), and a shift in the CBM occurs at a 5% decrease in $d$ for the SeMoS bilayer (Figure 6(d)). Figure S15 and Figure S16[SuppMat] present similar trends in the VBM and CBM ratios across all XMoY bilayers, further confirming the correlation between interlayer distance and band-edge contributions.
@@
As shown in Figure S15 (c) [SuppMat], a notable shift of the VBM from the K point to the $\Gamma$ point occurs in the MoTe2 bilayer with just a 1% decrease in $d$. This shift leads to a sudden increase in the VBM ratio, indicating a reduced depolarization effect from hole doping when $d$ is decreased beyond 1%. Interestingly, a similar behavior is observed for MoTe2, WSe2, and SWSe bilayers when SOC is turned off: the VBM shifts and the VBM ratio increase, as shown in Figure S8 [SuppMat], which results in a reduction in hole-doping-induced depolarization, as shown in Figure S9-S10 [SuppMat]. These observations suggest that decreasing the interlayer distance could likewise enhance the VBM ratio and reduce hole-doping-induced depolarization in WSe2 and SWSe bilayers.
@@
To validate our intuition regarding the enhancement of the VBM ratio with reduced interlayer distance ($d$) in WSe2 and SWSe bilayers, we analyze the variation of VBM and CBM ratios with $d$, as shown in Figure 6 (e)-(f). Both WSe2 and SWSe bilayers exhibit a sudden increase in the VBM ratio when $d$ is decreased beyond 3%, which corresponds to a shift of the VBM from the K point to the $\Gamma$ point, as shown in Figure 6(e). In contrast, the CBM ratio increases monotonically with increasing $d$, as shown in Figure 6 (f), similar to the behavior observed in the XMoY bilayers. The enhanced VBM ratios for WSe2 and SWSe bilayers closely resemble those obtained when the SOC effect is switched off, as shown in Figure S8 [SuppMat]. Therefore, we anticipate a comparable reduction in hole-doping-induced depolarization in WSe2 and SWSe bilayers for reduced $d$, consistent with the effect observed upon disabling SOC (Figure S10 [SuppMat]).
@@
Motivated by the potential for enhanced polarization with multiple stable polarization states [Deb2022, Cao2024], and the possibility of increased interlayer vdW interactions when only one Janus monolayer is present at the interface [Shengxi2020, Shengxi2021], we propose an ABC-stacked trilayer structure. In this design, a MoS2 monolayer is sandwiched between two Janus monolayers, with the intrinsic electric fields ($E$) in the Janus layers oriented toward the heterostructure interface, as illustrated in Figure 7(a).
The ABC stacking configuration shown in Figure 7(a) closely resembles the AB stacking illustrated in Figure 1(a), producing an upward out-of-plane polarization at each interface. The interlayer distances at the top and bottom interfaces, denoted as $d_{1}$ and $d_{2}$, respectively, are key factors in modulating the vdW interactions and the resulting interfacial polarizations. As shown in Table S4 [SuppMat], $d_{1}$ and $d_{2}$ have nearly identical values, indicating similar interlayer interactions and, consequently, comparable polarization at both interfaces, which is consistent with the behavior expected from interfacial ferroelectrics.
@@
The trilayer systems have lower (more negative) formation energies than their corresponding bilayer systems, indicating enhanced stability, as shown in Figure 7(b). We further validated the dynamical stability of both the Janus trilayers and their bilayer counterparts by calculating their phonon dispersion curves (Figure S18 [SuppMat]), which show no imaginary frequencies. Additionally, the thermal stability of the Janus trilayers and their bilayer counterparts was analyzed using AIMD simulations at 300 K and 600 K, as shown in the section Thermal Stability Analysis of the Supplementary Material [SuppMat]. No persistent structural changes were observed during the energy evolution, indicating robust thermal stability for these systems.
@@
In Figure 7(c), we compare the polarization values for the trilayer systems with those of their bilayer counterparts. Interestingly, similar to the bilayers, the trilayer Janus sliding ferroelectrics exhibit higher polarization when the $E$ fields point toward the interface. These $E$ fields also lead to a reduction in the interlayer distances, $d_{1}$ and $d_{2}$, as shown in Table S4 [SuppMat]. Additionally, the charge density profiles of the Janus trilayers, shown in Figure S19 [SuppMat], display similar modulation due to the $E$ fields as observed in the E-in stacked Janus bilayers. These results suggest that, in ABC-stacked Janus trialyers, the intrinsic $E$ fields modulate the polarization by redistributing interfacial electron density and tuning the interlayer spacing—mirroring the mechanism found in AB-stacked Janus bilayers.
The potential profiles in Figure S19 [SuppMat] confirm the interfacial nature of the polarization in the trilayer systems, which results in the cumulative nature of polarization at each interface. This explains the higher polarization values observed in Figure 7(c) compared to their bilayer counterparts. Additionally, the $\Delta\phi$ and $P_{CD}$ values follow a similar trend, further supporting the observed polarization behavior, as shown in Table S4 [SuppMat].
@@
The band structures of the trilayer systems exhibit strong asymmetry in the layer-wise contributions to the VBM and CBM. This leads to an uneven distribution of the partial charge density associated with the VBM and CBM, as illustrated in Figure S20 [SuppMat], with the asymmetry becoming more pronounced in the Janus sliding ferroelectric configurations. Consequently, based on the earlier discussion of doping-induced depolarization in Janus bilayers, we anticipate a significantly stronger depolarization effect from charge doping in trilayer systems compared to their bilayer counterparts.
@@
Finally, we would like to suggest some interesting future avenues for experimental and theoretical work motivated by our results. Experiments confirming our theoretical finding of enhancement of polarization along with bandgap and hole-doping-induced depolarization reduction by out-of-plane compressive strain for AB-stacked WSe2 bilayer would be an interesting and useful finding, considering the already existing experimental work on the doping-induced polarization modulation for WSe2 bilayer in Ref. [Deb2022]. Also, the semi-metallic nature observed after reducing the interlayer distance for the TeMoS bilayer suggests the possibility of designing Janus sliding ferroelectrics with intrinsic semi-metallicity by stacking them in E-in stacking, given much lower bandgaps for parent TMD bilayers. We expect this work to motivate several other theoretical and experimental investigations to design better-performing sliding ferroelectrics with additional properties such as a low bandgap and polar metallicity.
Supplementary Material
@@
Formation energy plot for bilayers; charge density and potential profiles for AB and BA stacking of SeMoS bilayer; charge density and potential profiles for XMoY bilayers; correlation plots for $\Delta\phi$ and $P_{CD}$ with $P_{berry}$ along with comparison of SOC and without SOC results; additional results for XMoY bilayers with $d=d_{avg}$; band structures of XMoY bilayers; comparison of band structure of MoTe2, WSe2 and SWSe bilayer without and with SOC; effect of doping on polarization for bilayers; variation of Fermi energy and DOS with doping for TeMoS and SMoTe bilayers; polarization and bandgap variation with more than 5% reduction in interlayer distance for XMoY bilayers along with band structures for bilayers showing semi-metallic nature at reduced interlayer distance; comparison of percentage change in bandgap and polarization with change in $d$ for XMoY bilayers; variation of the VBM and CBM ratios with change in $d$ for XMoY bilayers; variation of polarization and bandgap with change in $d$ for WSe2 and SWSe bilayers; phonon dispersion curves; charge density and potential profiles for trilayers; band structure with layer-wise contribution analysis for trilayers; tables for data for bilayer and trilayer systems presented and referred in the paper; methodology for $P_{CD}$ calculation; thermal stability analysis (PDF).